# LoongServe: Efficiently Serving Long-context Large Language Models with Elastic Sequence Parallelism

### v1

@@ Abstract.
The context window of large language models (LLMs) is rapidly increasing, leading to a huge variance in resource usage between different requests as well as between different phases of the same request. Restricted by static parallelism strategies, existing LLM serving systems cannot efficiently utilize the underlying resources to serve variable-length requests in different phases. To address this problem, we propose a new parallelism paradigm, elastic sequence parallelism (ESP), to elastically adapt to the variance between different requests and phases. Based on ESP, we design and build LoongServe, an LLM serving system that (1) improves computation efficiency by elastically adjusting the degree of parallelism in real-time, (2) improves communication efficiency by reducing key-value cache migration overhead and overlapping partial decoding communication with computation, and (3) improves GPU memory efficiency by reducing key-value cache fragmentation across instances. Our evaluation under diverse real-world datasets shows that LoongServe improves the maximum throughput by up to 3.85$\times$ compared to the chunked prefill and 5.81$\times$ compared to the prefill-decoding disaggregation.
@@ 1. Introduction
The emergence of large language models (LLMs) has fundamentally pushed modern applications into a new era. These LLMs are trained on vast data to gain extraordinary capabilities in a wide range of areas, such as programming (Nijkamp et al., 2023), chatting (cha, 2022) and planning (Gravitas, 2023). The context window is a key feature of LLMs. It is rapidly increasing to enable sophisticated reasoning about long documentation, relevant problem-solving with a large codebase, and customized generation based on long instructions (Google, 2024). Many organizations have released long-context LLMs, such as Anthropic’s Claude-3 (Anthropic, 2024), Google’s Gemini-1.5 (Google, 2024), and UC Berkeley’s Large World Model (LWM) (Liu et al., 2024), which all support a 1M context window.
@@
To accelerate the LLM computation, many works exploit different parallel dimensions. Model parallelism (Shoeybi et al., 2020) partitions the model parameters across multiple GPUs to parallelize the computation. Sequence parallelism (Korthikanti et al., 2022) partitions input sequences of requests across GPUs to achieve acceleration. Whether using one of them or a combination of both, existing practices decide the parallel configuration statically before launching the service.
@@
Experiments on real-world datasets show that LoongServe improves the maximum throughput by up to 3.85$\times$ compared to the chunked prefill and 5.81$\times$ compared to the prefill-decoding disaggregation. Furthermore, experiments show that LoongServe improves the performance of the prefill phase and decoding phase simultaneously.
@@
2. Background and Motivations
@@ 2.1. The Process of LLM Inference
Most of the popular LLMs (OpenAI, 2023; Anthropic, 2024; bar, 2023; Touvron et al., 2023a, b) adopt the Transformer architecture (Vaswani et al., 2017). The model consists of a stack of transformer layers, each containing an attention layer and a feed-forward network (FFN) layer. Attention layers make tokens in a request interact with other tokens, and FFN layers process requests token-wisely. In each iteration, given the tokens before it, the model predicts the next token. To avoid redundant computation, LLM serving systems cache the intermediate states of tokens, a.k.a Key-Value Cache, and reuse them for future token generation. This optimization divides the whole generation process into two phases: the prefill phase and the decoding phase. The prefill phase processes all the input tokens in a single iteration to build the key-value cache and generates the first output token, while the decoding phase only needs to compute the key-value cache for the newly generated output token. As a result, the prefill phase is much more compute-intensive than the decoding phase.
Figure 1. Sequence parallelism in the prefill phase.
@@ 2.2. Existing LLM Serving Systems
To accelerate LLM inference, existing solutions employ efficient GPU operator implementations, e.g., Flash Attention (Dao et al., 2022) and Flash Decoding (Dao, 2023) within a GPU, and exploit model parallelism, such as tensor parallelism (Shoeybi et al., 2020), to partition parameters of LLM across GPUs to parallelize the computation. However, the degree of model parallelism needs to be determined before launching the system.
To mitigate the impact of the long context, chunked prefill (Patel et al., 2023; Holmes et al., 2024) splits the long context into chunks and processes them chunks by chunks with the decoding phase, but still incurs interference between two phases (Zhong et al., 2024). To avoid interference, prefill-decoding disaggregation (Zhong et al., 2024) disaggregates two phases into different groups of GPUs. However, as mentioned in §1, it leads to high communication overhead due to frequent migration, GPU computation inefficiency due to resource mismatch, and GPU memory fragmentation due to isolation between groups.
@@ 2.3. Sequence Parallelism
To accelerate long-context LLM training, many LLM training systems propose sequence parallelism (Liu et al., 2023; Brandon et al., 2023; Jacobs et al., 2023; Li et al., 2023a; Korthikanti et al., 2023; Li et al., 2023c). In this work, we extend the Striped Attention (Brandon et al., 2023) to the LLM serving scenario.
Figure 1 illustrates its basic idea. Before processing, the input sequence is permuted, divided into multiple segments, and then dispatched to different instances. Except for the attention layer, instances execute without communication. At the attention layer, each instance simultaneously (1) computes the attention with its local query ($q_{i}$) and key-value ($kv_{j}$) tensors and (2) sends the key-value tensors ($kv_{j}$) to its neighbor instance ($instance_{(i+1)\%n}$). After multiple rounds of this process, query tensors in each instance interact with all key-value tensors and then generate the final output to the next layer. This sequence parallelism is compatible with popular attention mechanisms, such as Multi-Head Attention (MHA) (Vaswani et al., 2017), Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped-Query Attention (GQA) (Ainslie et al., 2023). It also has the same computational complexity as tensor parallelism.
However, it cannot directly apply to the LLM serving scenario. First, it only supports the prefill phase. Second, it is used in the LLM training scenario where the degree of parallelism in each training stage is fixed. Conversely, the LLM serving systems need to handle highly dynamic inference traffic and support unique characteristics introduced by the decoding phase, including unique computational patterns of the decoding phase and key-value cache management.
2.4. Motivations and Challenges
Despite existing LLM serving systems supporting long-context LLMs (§2.2), their static nature inherently mismatches with dynamic LLM serving workloads. LLM serving workloads is particularly dynamic in two aspects. First, as the context window of LLMs increases, the resource demand during the prefill phase varies significantly across requests with different input lengths in both computation and GPU memory consumption. As shown in Figure 2, processing 100K input tokens is 105.97 times slower than processing 1K input tokens. The difference in GPU memory consumption across requests can be as high as 1,000,000 times when serving an LLM with a 1M context window because the size of the key-value cache is linear to the input length. Secondly, even for a single request, the resource demand of the prefill and decoding phase differs vastly. As shown in Figure 2, because the prefill phase is compute-intensive, increasing DoP can significantly accelerate it. Conversely, for the relatively compute-lightweight decoding phase, a larger DoP may lead to negligible performance improvement due to additional communication overhead. Therefore, static parallelism strategies cannot be efficient in all scenarios.
@@ 2.4. Motivations and Challenges
Besides the inability to adapt to the dynamic workloads, they also cause GPU fragmentation issues. The reason is that the entire or most of the key-value cache of a request must reside in a single instance due to the locality constraint. This constraint leads to situations as depicted in Figure 4. Despite there being sufficient overall GPU memory (six slots), none of the instances can serve a request with six tokens due to the locality constraint.
To fundamentally address these issues, we propose elastic sequence parallelism (ESP). Our key insight is that LLM serving systems can extend SP to ESP by adding support to the decoding phase, flexibly managing the KV cache across phases, and dynamically distributing a request’s input tokens across instances, allowing for an adjustable degree of parallelism (DoP) without re-partitioning LLM parameters. As shown in Figure 3, using SP with existing parallelism strategies, such as tensor parallelism, does not introduce additional overheads and may even achieve better performance for requests with different lengths in both context and decoding phases, indicating a potential to further LLM inference acceleration. Moreover, the Elastic Sequence Parallelism (ESP) facilitates the dynamic utilization of memory resources across multiple instances, thereby mitigating the problem of memory fragmentation.
@@
While the dynamic adjustment capability of Elastic Sequence Parallelism (ESP) to meet the computational demands of diverse requests is indeed promising, unleashing its full potential presents several challenges.
@@ 3. LoongServe Overview
To address these problems, we propose a new parallelism strategy, Elastic Sequence Parallelism (ESP), and build a distributed LLM serving system, LoongServe, to fully unleash the potential of ESP. Figure 5 shows the architecture of LoongServe. LoongServe consists of a set of elastic instances and a global manager. These elastic instances can dynamically organize themselves into a set of disjoint ESP groups to process batches of requests in parallel with different configurable degrees of parallelism. They can also support efficient elastic scaling up and down without additional overhead to adapt to the varying resource demands of requests (§4). The GPU memory of LoongServe elastic instances also forms a unified distributed key-value cache pool, which can be flexibly used to store key-value tensors of requests at the granularity of a token across elastic instances to reduce GPU memory fragmentation (§4). The LoongServe global manager is responsible for managing requests, elastic instances, and the unified distributed key-value cache pool with the global view.
In each iteration, based on profiling results from the scaling information base (SIB), the global manager dynamically adjusts the degree of parallelism of existing batches, grouping strategy of elastic instances, batching and dispatching strategy of newly arrived requests, and placement strategy of key-value caches to improve the throughput and reduce the latency of requests in real-time (§5). In this process, the LoongServe dispatcher accompanied by the global manager dispatches newly arrived requests to a set of specific elastic instances as the global manager requires. At the same time, given the DoP and scaling plan generated by the global manager, the elasticity controller orders elastic instances to update their configuration to form the corresponding ESP groups and process requests in parallel. The key-value caches generated from LLM inference are placed at the position as specified in the scaling plan. The global manager monitors the progress of requests, the resource usage of elastic instances, and the key-value cache pool to update its decision.
@@ 4. LoongServe Elastic Instances
Elastic instances are the minimum independent execution units in LoongServe. Each maintains a replica of the model weights and employs a unified strategy of model parallelisms deployed on the equivalent number of GPUs (§5.4). Before each iteration, the global manager dynamically assigns elastic instances into multiple parallel groups, where each group handles the computation for a specific batch with different degrees of parallelism (DoP). The number of instances in a parallel group is its corresponding degree of sequence parallelism (DoSP). Elastic instances may also receive potential elastic scaling plans from the global manager, which entails the reassignment of instances’ parallel groups. If so, upon completion of the iteration’s computation, instances are required to realize elastic scaling to ensure that subsequent computations can proceed according to the newly defined DoP, meaning that the necessary key-value tensors are already in each instances’s KV Cache pool.
@@
Figure 6 shows the lifecycle of requests, where $(B_{i},I_{j})$ indicates the $j$-th iteration of batch $i$. Because the computational complexity of the prefill phase is much higher than that of the decoding phase, it is always necessary to scale down the batch after the prefill phase ($(B_{1},I_{1})$ and $(B_{2},I_{1})$). As the generated tokens become more and more, the computational complexity of the decoding phase increases and the GPU memory in a parallel group may be filled up with the KV Cache, leading to the necessity of scaling up the parallel group ($(B_{1},I_{3})$ and $(B_{2},I_{2})$). It can also optionally scale down the decoding batch to leave more resources for the prefill batch ($(B_{2},I_{3})$).
To mitigate the overhead of scaling mechanisms, we have devised a set of zero-overhead ESP mechanisms, which allow for elastic scaling down in the prefill phase and scaling up in the decoding phase without any additional key-value tensor migrations. In this section, we elaborate on the design of these mechanisms. As for the optional scaling down in the decoding phase, the global manager only uses it when its benefits outweigh the overhead of scaling down (§5).
@@ 4.1. Elastic Scale-down
Existing solution: reactive migration. Existing practices (Zhong et al., 2024) are reactive migration that migrates the key-value tensors from the parallel group $R$ to the new parallel group $R^{\prime}$ after the prefill phase. However, the migration overhead is non-negligible and escalates linearly with the sequence length of requests. Specifically, for a request with an input length of 500K, the GPU consumption of the key-value tensors exceeds 2.2TB when serving a 175B LLM. Even if GPUs are connected by high-bandwidth interconnects, such as NVLINK and Infiniband, migrating a single request still spans several seconds, significantly longer than a decoding step.
@@ 4.2. Elastic Scale-up
Existing solution: entire requests migration. Previous works such as tensor-parallelism (Shoeybi et al., 2020) (TP) and FlashDecoding (Dao, 2023) only support distributed decoding computations across multiple GPUs within a single instance. When resources of an instance, such as GPU memory, are insufficient, they have to migrate some requests in the batch to another instance, and then parallelly process them in different instances. However, migrating entire key-value caches of a request to another instance incurs significant migration overhead, even higher than a decoding step itself. Additionally, it requires the entire or most of the key-value tensors of a request must be stored in a single instance, leading to memory fragmentation issues.

### v2

@@ Abstract.
The context window of large language models (LLMs) is rapidly increasing, leading to a huge variance in resource usage between different requests as well as between different phases of the same request. Restricted by static parallelism strategies, existing LLM serving systems cannot efficiently utilize the underlying resources to serve variable-length requests in different phases. To address this problem, we propose a new parallelism paradigm, elastic sequence parallelism (ESP), to elastically adapt to the variance across different requests and phases. Based on ESP, we design and build LoongServe, an LLM serving system that (1) improves computation efficiency by elastically adjusting the degree of parallelism in real-time, (2) improves communication efficiency by reducing key-value cache migration overhead and overlapping partial decoding communication with computation, and (3) improves GPU memory efficiency by reducing key-value cache fragmentation across instances. Our evaluation under diverse real-world datasets shows that LoongServe improves the throughput by up to 3.85$\times$ compared to chunked prefill and 5.81$\times$ compared to prefill-decoding disaggregation.
Inference Serving; Large Language Models; Elastic Sequence Parallelism
ccs: Computer systems organization Real-time system architectureccs: Computing methodologies Natural language processingccs: Computing methodologies Distributed computing methodologiesjournalyear: 2024copyright: acmlicensedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695948isbn: 979-8-4007-1251-7/24/11
@@ 1. Introduction
The emergence of large language models (LLMs) has fundamentally pushed modern applications into a new era. These LLMs are trained on vast data to gain extraordinary capabilities in a wide range of areas, such as programming (Nijkamp et al., 2023), chatting (cha, 2022) and planning (Gravitas, 2023). The context window is a key feature of LLMs. It is rapidly increasing to enable sophisticated reasoning about long documentation, relevant problem-solving with a large codebase, and customized generation based on long instructions (Google, 2024). Many organizations have released long-context LLMs, such as Anthropic’s Claude-3 (Anthropic, 2024), Google’s Gemini-1.5 (Google, 2024), and UC Berkeley’s Large World Model (LWM) (Liu et al., 2024b), which all support a 1M context window.
@@
To accelerate the LLM computation, many works exploit different parallel dimensions. Model parallelism (Shoeybi et al., 2019) partitions the model parameters across multiple GPUs to parallelize the computation. Sequence parallelism (Korthikanti et al., 2022) partitions input sequences of requests across GPUs to achieve acceleration. Whether using one of them or a combination of both, existing practices decide the parallel configuration statically before launching the service.
@@
Experiments on real-world datasets show that LoongServe improves the throughput by up to 3.85$\times$ compared to chunked prefill and 5.81$\times$ compared to prefill-decoding disaggregation. Furthermore, experiments show that LoongServe improves the performance of the prefill phase and decoding phase simultaneously.
@@
2. Background and Motivation
Figure 1. Sequence parallelism in the prefill phase.
@@ 2.1. The Process of LLM Inference
Most of the popular LLMs (Achiam et al., 2023; Anthropic, 2024; bar, 2023; Touvron et al., 2023a, b) adopt the Transformer architecture (Vaswani et al., 2017). A model typically consists of a stack of transformer layers, each containing an attention layer and a feed-forward network (FFN) layer. Attention layers make tokens in a request interact with other tokens, and FFN layers process requests token-wisely. In each iteration, given the tokens before it, the model predicts the next token. To avoid redundant computation, LLM serving systems cache the intermediate states of tokens, a.k.a key-value cache, and reuse them for future token generation. This optimization divides the whole generation process into two phases: the prefill phase and the decoding phase. The prefill phase processes all the input tokens in a single iteration to build the key-value cache and generates the first output token, while the decoding phase only needs to compute the key-value cache for the newly generated output token. As a result, the prefill phase is more compute-intensive than the decoding phase.
@@ 2.2. Existing LLM Serving Systems
To accelerate LLM inference, existing solutions employ efficient GPU kernel implementations, e.g., Flash Attention (Dao et al., 2022) and Flash Decoding (Dao, 2023) within a GPU, and exploit model parallelism, such as tensor parallelism (Shoeybi et al., 2019), to partition model parameters across GPUs to parallelize the computation. However, the degree of model parallelism needs to be determined before launching the system.
To mitigate the impact of the long context, chunked prefill (Patel et al., 2023; Holmes et al., 2024) splits the long context into chunks and processes them chunk by chunk with the decoding phase, but still incurs interference between two phases (Zhong et al., 2024). To avoid interference, prefill-decoding disaggregation (Zhong et al., 2024) disaggregates two phases into different groups of GPUs. However, as mentioned in §1, it leads to high communication overhead due to frequent migration, GPU computation inefficiency due to resource mismatch, and GPU memory fragmentation due to isolation between groups.
@@ 2.3. Sequence Parallelism
To accelerate long-context LLM training, many LLM training systems propose sequence parallelism (Liu et al., 2023; Brandon et al., 2023; Jacobs et al., 2023; Li et al., 2023a; Korthikanti et al., 2023; Li et al., 2023c). In this work, we extend Striped Attention (Brandon et al., 2023) to the LLM serving scenario.
@@
Figure 1 illustrates its basic idea. Before processing, the input sequence is permuted, divided into multiple segments, and then dispatched to different instances. Except for the attention layer, the instances execute without communication. At the attention layer, each instance simultaneously (1) computes the attention with its local query ($q_{i}$) and key-value ($kv_{j}$) tensors and (2) sends the key-value tensors ($kv_{j}$) to its neighbor instance ($instance_{(i+1)\%n}$). After multiple rounds of this process, query tensors in each instance interact with all key-value tensors and then generate the final output to the next layer. Sequence parallelism is compatible with popular attention mechanisms, such as Multi-Head Attention (MHA) (Vaswani et al., 2017), Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped-Query Attention (GQA) (Ainslie et al., 2023). It also has the same computational complexity as tensor parallelism and consumes less GPU memory for buffer activations (Liu et al., 2023). Furthermore, it can be used in conjunction with other parallelisms, e.g., tensor parallelism, to accelerate LLM training further.
However, it cannot be directly applied to the LLM serving scenario. First, it only supports the prefill phase. Second, it is used in the LLM training scenario where the degree of parallelism in each training stage is fixed. Conversely, the LLM serving systems need to handle highly dynamic inference traffic and support unique characteristics introduced by the decoding phase, including unique computational patterns of the decoding phase and key-value cache management.
2.4. Motivation and Challenges
Despite existing LLM serving systems supporting long-context LLMs (§2.2), there is an inherent mismatch between their static nature and the dynamics in LLM serving. Particularly, LLM serving workloads are dynamic in two aspects. First, as the context window of LLMs increases, the resource demand during the prefill phase varies significantly across requests with different input lengths in both computation and GPU memory consumption. As shown in Figure 2, processing 100K input tokens for a LWM model on 8 GPUs is 105.97 times slower than processing 1K input tokens. The difference in GPU memory consumption across requests can be as high as 1,000,000 times when serving an LLM with a 1M context window because the size of the key-value cache is linear to the input length. Secondly, even for a single request, the resource demand of the prefill and decoding phase differs vastly. As shown in Figure 2, because the prefill phase is compute-intensive, increasing DoP can significantly accelerate it. Conversely, for the relatively compute-lightweight decoding phase, a larger DoP may lead to negligible performance improvement due to additional communication overhead. Therefore, static parallelism strategies cannot be efficient in all scenarios.
@@ 2.4. Motivation and Challenges
Besides the inability to adapt to the dynamics, existing solutions (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2023; Zhong et al., 2024; Hu et al., 2024) also cause GPU fragmentation issues. The reason is that the entire or most of the key-value cache of a request must reside in a single instance due to the locality constraint. This constraint leads to situations as depicted in Figure 4. Despite there being sufficient overall GPU memory (six slots), none of the instances can serve a request with six tokens due to the locality constraint.
Elastic sequence parallelism. To fundamentally address these issues, we propose elastic sequence parallelism (ESP). Our key insight is that LLM serving systems can extend SP to ESP by adding support to the decoding phase, flexibly managing the KV cache across phases, and dynamically distributing a request’s input tokens across instances, allowing for an adjustable DoP without re-partitioning LLM parameters. As shown in Figure 3, using SP with existing parallelism strategies, such as tensor parallelism, does not introduce additional overheads and may even achieve better performance for requests with different lengths in both context and decoding phases, indicating a potential to further LLM inference acceleration. Moreover, ESP facilitates flexible utilization of memory resources across multiple instances, thereby mitigating the problem of memory fragmentation.
@@
Challenges. While the dynamic adjustment capability of ESP to meet the computational demands of diverse requests is indeed promising, unleashing its full potential presents several challenges.
@@ 3. LoongServe Overview
To address these challenges, we design and build a distributed LLM serving system, LoongServe, to fully unleash the potential of ESP. Figure 5 shows the architecture of LoongServe. LoongServe consists of a set of elastic instances and a global manager. These elastic instances can dynamically organize themselves into a set of disjoint ESP groups to process batches of requests in parallel with different configurable degrees of parallelism. They can also support efficient elastic scaling up and down without additional overhead to adapt to the varying resource demands of requests (§4). The GPU memory of LoongServe elastic instances also forms a unified distributed key-value cache pool, which can be flexibly used to store key-value tensors of requests at the granularity of a token across elastic instances to reduce GPU memory fragmentation (§4). The LoongServe global manager is responsible for managing requests, elastic instances, and the unified distributed key-value cache pool with the global view.
In each iteration, based on profiling results from the scaling information base (SIB), the global manager dynamically adjusts the DoP of existing batches, grouping strategy of elastic instances, batching and dispatching strategy of newly arrived requests, and placement strategy of key-value caches to improve the throughput and reduce the latency of requests in real-time (§5). In this process, the LoongServe dispatcher accompanied by the global manager dispatches newly arrived requests to a set of specific elastic instances as the global manager requires. At the same time, given the DoP and scaling plan generated by the global manager, the elasticity controller orders elastic instances to update their configuration to form the corresponding ESP groups and process requests in parallel. The key-value caches generated from LLM inference are placed at the position as specified in the scaling plan. The global manager monitors the progress of requests, the resource usage of elastic instances, and the key-value cache pool to update its decision.
@@ 4. LoongServe Elastic Instances
Elastic instances are the minimum independent execution units in LoongServe. Each maintains a replica of the model weights and employs a unified model parallelisms on the equivalent number of GPUs (§5.4). Before each iteration, the global manager dynamically assigns elastic instances into multiple parallel groups, where each group handles the computation for a specific batch with different degrees of parallelism (DoP). The number of instances in a parallel group affects its corresponding DoP. Elastic instances may also receive potential elastic scaling plans from the global manager, which entails the reassignment of instances’ parallel groups. If so, upon completion of the iteration’s computation, instances are required to realize elastic scaling to ensure that subsequent computations can proceed according to the newly defined DoP, meaning that the necessary key-value tensors are already in each instances’s KV Cache pool.
@@
Figure 6 shows the lifecycle of requests, where $(B_{i},I_{j})$ indicates the $j$-th iteration of batch $i$. Because the computational complexity of the prefill phase is much higher than that of the decoding phase, it is always necessary to scale down the batch after the prefill phase ($(B_{1},I_{1})$ and $(B_{2},I_{1})$). As the generated tokens become more and more, the computational complexity of the decoding phase increases and the GPU memory in a parallel group may be filled up with the key-value cache, leading to the necessity of scaling up the parallel group ($(B_{1},I_{3})$ and $(B_{2},I_{2})$). It can also optionally scale down the decoding batch to leave more resources for the prefill batch ($(B_{2},I_{3})$).
To mitigate the overhead of scaling mechanisms, we have design a set of zero-overhead ESP mechanisms, which allow for elastic scaling down in the prefill phase and scaling up in the decoding phase without any additional key-value tensor migrations. In this section, we elaborate on the design of these mechanisms. As for the optional scaling down in the decoding phase, the global manager only uses it when its benefits outweigh the overhead of scaling down (§5).
@@ 4.1. Elastic Scale-down
Existing solution: reactive migration. Existing practices (Patel et al., 2023; Zhong et al., 2024; Hu et al., 2024) use reactive migration that migrates the key-value tensors from the parallel group $R$ to the new parallel group $R^{\prime}$ after the prefill phase. However, the migration overhead is non-negligible and escalates linearly with the sequence length of requests. Specifically, for a request with an input length of 1M, the GPU consumption of the key-value tensors exceeds 488GB when serving a 7B LLM. Even if GPUs are connected by high-bandwidth interconnects, such as NVLINK and Infiniband, migrating a single request still spans several seconds, significantly longer than a decoding step.
@@ 4.2. Elastic Scale-up
Existing solution: entire requests migration. Previous works such as tensor-parallelism (Shoeybi et al., 2019) (TP) and FlashDecoding (Dao, 2023) only support distributed decoding computations across multiple GPUs within a single instance. When resources of an instance, such as GPU memory, are insufficient, they have to migrate some requests in the batch to another instance, and then parallelly process them in different instances. However, migrating entire key-value caches of a request to another instance incurs significant migration overhead, even higher than a decoding step itself. Additionally, it requires the entire or most of the key-value tensors of a request must be stored in a single instance, leading to memory fragmentation issues.